# Large Language Model Sentinel: LLM Agent for Adversarial Purification

### v3

@@ Methods
      FS Standard w/o LLAMOS Character Word Sentence
@@
      ✓ 92.18 30.42 85.16 79.13 68.96
Related Work
Adversarial Attack. Deep neural networks (DNNs) are vulnerable to adversarial examples (Szegedy et al. 2014), which are generated by adding small, human-imperceptible perturbations to natural examples, but completely change the prediction results to DNNs (Goodfellow, Shlens, and Szegedy 2015; Lin et al. 2024b). With the rapidly increasing applications of LLMs (OpenAI 2023; Köpf et al. 2024; Romera-Paredes et al. 2024), security concerns have emerged as a critical area of research (Gehman et al. 2020; Bender et al. 2021; Mckenna et al. 2023; Manakul, Liusie, and Gales 2023; Liu et al. 2023b; Zhu et al. 2023; Li, Song, and Qiu 2023; Qi et al. 2024; Yao et al. 2024b), with researchers increasingly focusing on adversarial attacks targeting LLMs. In a similar setup to DNNs, for LLMs, attackers manipulate a small amount of text to change the output of the target LLM while maintaining the semantic information for humans (Wang et al. 2024; Xu et al. 2024). Presently, addressing the security issues surrounding LLMs is of paramount importance and requires urgent attention.
Adversarial Defense. There are two main defense techniques on traditional DNNs, including adversarial training (AT) (Goodfellow, Shlens, and Szegedy 2015) and adversarial purification (AP) (Shi, Holtz, and Mishne 2021; Srinivasan et al. 2021). Unlike traditional DNNs, retraining LLMs is nearly impossible due to cost issues (Li, Song, and Qiu 2023). Therefore, most methods enhance the robustness of LLMs through adversarial fine-tuning (AFT) (Xiang et al. 2024; Li et al. 2024b; Bianchi et al. 2023; Deng et al. 2024; Qi et al. 2024). While AFT can effectively defend against attacks, it remains susceptible to unseen attacks whose adversarial examples that the LLMs have not previously learned (Li, Song, and Qiu 2023). Additionally, even with fine-tuning, training the LLMs will still consume a significant cost (Hu et al. 2022; Dettmers et al. 2024). Adversarial purification (AP) aims to purify adversarial examples before feeding them into the target model, which has emerged as a promising defense method (Shi, Holtz, and Mishne 2021; Srinivasan et al. 2021; Lin et al. 2024b). Compared with the AT or AFT method, the AP method utilizes an additional model that can defend against unseen attacks without retraining the target model (Lin et al. 2024b; Li, Song, and Qiu 2023). In some traditional computer vision and natural language processing tasks, researchers have started using LLMs as purifiers for adversarial purification (Singh and Subramanyam 2024; Li et al. 2024a; Moraffah et al. 2024), but the security issues of LLMs themselves have not been deeply considered. Therefore, we propose a novel LLM defense technique named LLAMOS to purify the adversarial textual examples before feeding them into the target LLM, aiming to improve the robustness.
Large Language Model Agent. The LLM agent is a new research direction that has emerged in recent years (Ha, Florence, and Song 2023; Mu et al. 2024; M. Bran et al. 2024). This novel type of agent is capable of interacting with humans in natural language, leading to a significant increase in applications across fields such as chatbots, natural sciences, robotics, and workflows (Boiko, MacKnight, and Gomes 2023; Yang et al. 2023; Lin et al. 2024a; Wang et al. 2023b; Liu et al. 2023a). Furthermore, LLMs have demonstrated promising zero-shot/few-shot planning and reasoning capabilities across various configurations (Sumers et al. 2023), covering specific environments and reasoning tasks (Yao et al. 2023; Gong et al. 2023; Yao et al. 2024a). In this paper, we introduce a new variant of the LLM agent designed specifically to purify adversarial textual examples generated by attacks.
Experiments
In this section, we conduct extensive experiments on GLUE datasets to evaluate the effectiveness of the proposed method (LLAMOS). Specifically, our method significantly reduces the attack success rate (ASR) by up to 37.86% with GPT-3.5 and 45.59% with LLAMA-2, respectively.
Experimental Setup
Datasets. The experiments are conducted on six tasks in GLUE datasets (Wang et al. 2018), including SST-2, RTE, QQP, QNLI, MNLI-mm, MNLI-m (Socher et al. 2013; Dagan, Glickman, and Magnini 2005; Bar-Haim et al. 2006; Giampiccolo et al. 2007; Bos and Markert 2005; Bentivogli et al. 2009; Wang, Hamza, and Florian 2017; Rajpurkar et al. 2016; Williams, Nangia, and Bowman 2018). The detailed descriptions are provided in Appendix A.
Adversarial Attacks. We evaluate our method against PromptAttack (Xu et al. 2024), which is a powerful attack that combines nine different types of attacks, as illustrated in Table 11. Furthermore, Xu et al. (2024) introduce the few-shot (FS) strategy (Logan IV et al. 2021) and ensemble (EN) strategy (Croce and Hein 2020) to boost the attack power of PromptAttack, details in Appendix B.
Evaluation Metrics. We evaluate the performance of defense methods using two metrics: attack success rate (ASR) and robust accuracy (RA). These metrics are derived from testing on adversarial examples, where a lower ASR or a higher RA indicates greater model robustness.
Training Details. The experiments in this paper are conducted using GPT-3.5 (OpenAI 2023) with ‘GPT-3.5-Turbo-0613’ version and LLAMA-2 (Touvron et al. 2023b) with ‘LLAMA-2-7b’ version. For GPT-3.5, we purchase OpenAI’s API service [^2] https://openai.com/api/ and conduct testing experiments with the ‘openai’ package in Python. For LLAMA-2, we deploy it locally on NVIDIA RTX A6000 and utilize the available checkpoint published by MetaAI from HuggingFace [^3] https://huggingface.co/meta-llama/.
Table 5: Standard accuracy and robust accuracy defense against PromptAttack with GPT-3.5. In the first three lines, we present the average robust accuracy across different attacks to more clearly highlight the improvements of LLAMOS. The remaining lines show the robust accuracy of LLAMOS to defend against various attacks.
@@
Results
Evaluation of LLAMOS Performance on Attack Success Rate (ASR). We evaluate the ASR of the LLAMOS against PromptAttack-EN and PromptAttack-FS-EN on the GLUE datasets with GPT-3.5 (OpenAI 2023). As shown in Table 2, our method significantly reduces the ASR of both PromptAttack-EN and PromptAttack-FS-EN across all tasks. Specifically, our method achieves an average ASR reduction of 29.33% and 29.39%, respectively. These results demonstrate that LLAMOS is effective in defending against adversarial textual attacks. Additionally, we also evaluate the performance of LLAMOS on the SST-2 dataset with LLAMA-2 (Touvron et al. 2023b), as shown in Table 3. The results are similar to the previous experiments. The ASR of PromptAttack-EN and PromptAttack-FS-EN is significantly reduced by 45.59% and 10.57%, respectively.
Evaluation of LLAMOS Performance on Robust Accuracy (RA). We evaluate the RA of LLAMOS on the SST-2 dataset with LLAMA-2 against three types of PromptAttack: character, word, and sentence attacks. In Table 4, the accuracies in the first two columns represent the standard accuracy and robust accuracy without defense, while the last three columns represent the robust accuracy with LLAMOS. Under strong attacks, the classification accuracy of the target LLM decreased from 92.18% to 30.42%. LLAMOS can effectively defend against adversarial textual attacks, significantly improving the robust accuracy. Specifically, the lowest robust accuracy reaches 86.96%. Additionally, we conduct comprehensive experiments across nine types of attacks and six tasks with GPT-3.5, as shown in Table 5. LLAMOS can effectively defend against character-level attacks, achieving results on C1 and C3 that closely match the standard accuracy.
Table 6: Comparison on robust accuracy against C3 attack with in-context learning.
       ICL C3 C3 with few-shot
@@ Results
       ✓ 97.66 92.19
Table 7: Robust accuracy defense against the attack agent under different iterations with GPT-3.5.
      Iterations Round 1 Round 2 Round 3 Round 4 Round 5
@@
Table 8: Comparison between clean examples, adversarial examples, and purified examples. When incorrectly classified, the modified area is highlighted with a yellow background.
@@
       LLAMOS+ICL Purified: a fast , funny , highly enjoyable movie .. Positive
The purified and adversarial examples generated by the agents in the adversarial system during iterations.
@@
Adversarial: Youdo n’thave to know music to appreciate the film’s easygoing … Negative
       Round 2 Purified: You don’t have to know music to appreciate the film’s easygoing … Positive
Adversarial: You don’t have to know music todepreciatethe film’s easygoing … Negative
       Round 3+ Purified: You don’t have to know music to appreciate the film’s easygoing … Positive
Adversarial: You don’t have to know music todepreciatethe film’s easygoing … Negative
@@
Evaluation of LLAMOS Performance with In-Context Learning (ICL). The C3-based attack (Xu et al. 2024) is a very obvious attack that adds up to two extraneous characters to the end of the sentence, as shown in Table 8. However, our method only achieves robust accuracies of 89.06% for the C3 attack and 66.41% for the C3-FS attack, respectively. To further improve the robustness, we introduce ICL to enhance the performance of the defense agent. As shown in Table 6, the defense agent with ICL significantly improves the robust accuracy against the C3 attack, achieving a robust accuracies of 97.66% for the C3 attack and 92.19% for the C3-FS attack.
@@
Analysis of Adversarial System. We conduct experiments with an adversarial system and evaluate the robust accuracy defense against adversarial examples generated by the attack agent over multiple iterations. As shown in Table 7, the defense agent initially achieves a robust accuracy of 96.09% in the first round of confrontation. However, after the purified examples are re-attacked by the attack agent, the robust accuracy decreases to 56.25%. The defense agent then purifies these adversarial examples again, leading to an increase in robust accuracy, but it will decrease once more by subsequent attacks. This continual fluctuation in robust accuracy is a common phenomenon in adversarial training (Goodfellow, Shlens, and Szegedy 2015). Upon reviewing the generated texts, we observe that after several rounds of confrontation, both the defense agent and attack agent may generate the same sentences as previous ones, resulting in a potential infinite loop, as shown in Table 8. This is an interesting phenomenon that requires further investigation, particularly strategies to disrupt such loops.
Discussion
@@
The Advantages of LLAMOS. As emphasized by the experimental results presented in the Results Section and Table 12, LLAMOS significantly enhances performance across various tasks and attacks with LLAMA-2 and GPT-3.5. Additionally, the defense agent in LLAMOS is a plug-and-play module, serving as a pre-processing step. Through in-context learning (ICL), the defense agent can be continuously optimized to defend against emerging attacks. This invisibly resolves a major challenge in adversarial robustness: Due to the significant differences between different attacks, the model trained on specific attacks often fails to generalize to other unseen attacks (Poursaeed et al. 2021; Laidlaw, Singla, and Feizi 2021; Tack et al. 2022). The model necessitates to be continuously fine-tuning to adapt to emerging attacks. However, fine-tuning parameters requires substantial costs (Hu et al. 2022; Dettmers et al. 2024), and the emergence of new attack techniques continually makes it impractical to train the model to defend against emerging attacks (Laidlaw, Singla, and Feizi 2021; Lin et al. 2024b). In contrast, our method can effectively enhance the robustness through ICL without adjusting the parameters of the LLMs, which is undoubtedly a significant advantage.
@@
The Challenges in LLM-based Defense. The defense agent is tasked with purifying adversarial examples, but it is difficult to distinguish between natural examples and adversarial examples in some cases. As shown in Table 12.4, the attacker altered the original meaning by inserting ‘not’, rendering the adversarial example indistinguishable from a natural example, resulting in the defense agent failing to generate the correct sentence. Although we hope that the defense agent can observe the sentences like humans, it presents a huge challenge at present. Unlike the attacker or humans, the defense agent lacks access to the original label of the input sentence. Furthermore, although the defense agent can effectively defend against adversarial attacks, it cannot prevent subsequent attacks, as illustrated in Table 7. For instance, the malicious LLMs can embed specific system prompts to influence the output, which is unbeknownst to users; they can add ‘:)’ to each input sentence for prediction rather than predicting the original sentence. In this case, the defense agent also fails to defend. This issue is also an important problem in traditional adversarial training (Goodfellow, Shlens, and Szegedy 2015), and no method has completely resolved this issue. Nonetheless, as previously discussed, LLMs offer advantages not available to traditional DNNs, and we have naturally solved one challenge in adversarial robustness, which is that the model can adapt to new attacks. Hence, future advancements may resolve adversarial issues of attack and defense within LLM frameworks, representing a challenging but promising research direction.
Conclusion
@@
In this paper, we propose LLAMOS, a novel LLM-based defense technique designed to purify adversarial examples before feeding them into the target LLM. The defense agent within LLAMOS operates as a plug-and-play module that functions effectively as a pre-processing step without requiring retraining of the target LLM. We conduct extensive experiments across various tasks and attacks with LLAMA-2 and GPT-3.5. The results demonstrate that LLAMOS can effectively defend against adversarial attacks. Furthermore, we discuss certain existing shortcomings and challenges, which we aim to address in future research. Ultimately, we discuss the Limitations and Impact Statements in the Appendix D.

### v4

@@ 5.1.4 Experiment Details
FS SA RA Character Word Sentence
@@
92.18 30.42 85.16 79.13 68.96
Table 5: Standard accuracy and robust accuracy of GPT-3.5 against PromptAttack. The first three rows report average accuracy across all attacks; the remaining rows give robust accuracy for each individual attack.
@@
5.2 Results
Performance on ASR: We evaluate the ASR against PromptAttack-EN and PromptAttack-FS-EN on the GLUE datasets with GPT-3.5 [22]. As shown in Table 2, our method significantly reduces the ASR of both PromptAttack-EN and PromptAttack-FS-EN across all tasks. Specifically, our method achieves an average ASR reduction of 29.33% and 29.39%, respectively. These results demonstrate that LLAMOS is effective in defending against adversarial textual attacks. Additionally, we also evaluate the performance of LLAMOS on the SST-2 dataset with LLAMA-2 [32], as shown in Table 3. The results are similar to the previous experiments. The ASR of PromptAttack-EN and PromptAttack-FS-EN is significantly reduced by 45.59% and 10.57%, respectively.
Performance on RA: We evaluate the RA on the SST-2 dataset with LLAMA-2 against three types of PromptAttack: character, word, and sentence attacks. In Table 4, the first two columns represent the standard accuracy (SA) and robust accuracy (RA) without defense, while the last three columns represent the robust accuracy with LLAMOS. Under strong attacks, the accuracy of the target LLM decreased from 92.18% to 30.42%. LLAMOS can effectively defend against adversarial textual attacks, significantly improving the robust accuracy. Specifically, the lowest robust accuracy reaches 86.96%. Additionally, we conduct comprehensive experiments across nine types of attacks and six tasks with GPT-3.5, as shown in Table 5. LLAMOS can effectively defend against character-level attacks, achieving results on C1 and C3 that closely match the standard accuracy.
Table 6: Comparison on robust accuracy against C3 attack with ICL.
ICL C3 C3 with FS
@@ 5.2 Results
97.66 92.19
Table 7: Robust accuracy of agents over successive rounds (R.) in the adversarial system. Each column shows accuracy after each attack–defense iteration.
Iterations R. 1 R. 2 R. 3 R. 4 R. 5
@@
Table 8: Comparison between clean, adversarial, and purified examples. When incorrectly classified, the modified area is highlighted with a yellow background.
@@
+ICL Purified: a fast , funny , highly enjoyable movie .. Positive
The purified and adversarial examples generated in the adversarial system.
@@
Adversarial: Youdo n’thave to know music to appreciate … Negative
Round 2 Purified: You don’t have to know music to appreciate … Positive
Adversarial: You don’t have to know music todepreciate Negative
Round 3 Purified: You don’t have to know music to appreciate … Positive
Adversarial: You don’t have to know music todepreciate Negative
@@
LLAMOS with In-Context Learning: The C3 attack [41] is a very obvious attack that adds up to two extraneous characters to the end of the sentence, as shown in Table 8. However, our method only achieves robust accuracies of 89.06% for the C3 attack and 66.41% for the C3-FS attack. To further improve the robustness, we introduce ICL to enhance the performance of the defense agent. As shown in Section 5.2, the defense agent with ICL significantly improves the robust accuracy against the C3 attack, achieving a robust accuracies of 97.66% for the C3 attack and 92.19% for the C3-FS attack.
@@
Analysis of Adversarial System: We conduct experiments with an adversarial system and evaluate the robust accuracy against adversarial examples generated by the attack agent over multiple iterations. As shown in Table 7, the defense agent initially achieves a robust accuracy of 96.09% in the first round of confrontation. However, after the purified examples are re-attacked by the attack agent, the robust accuracy decreases to 56.25%. The defense agent then purifies these adversarial examples again, leading to an increase in robust accuracy, but it will decrease once more by subsequent attacks. This continual fluctuation in robust accuracy is a common phenomenon in adversarial training [5]. Upon reviewing the generated texts, we observe that after several rounds of confrontation, both the defense agent and attack agent may generate the same sentences as previous ones, resulting in a potential infinite loop, as shown in Table 8. This is an interesting phenomenon that requires further investigation, particularly strategies to disrupt such loops.
5.3 Discussion
5.3.1 The Advantages of LLAMOS
@@
As emphasized by the experimental results presented in the Results Section and Table 12, LLAMOS significantly enhances performance across various tasks and attacks with LLAMA-2 and GPT-3.5. Additionally, the defense agent in LLAMOS is a plug-and-play module, serving as a pre-processing step. Through ICL, the defense agent can be continuously optimized to defend against emerging attacks. This invisibly resolves a major challenge in adversarial robustness: Due to the significant differences between different attacks, the model trained on specific attacks often fails to generalize to other unseen attacks [8]. As a result, continuous fine-tuning is needed to stay robust against emerging attacks, but this incurs substantial costs [6, 3], and the emergence of attack techniques continually makes it impractical to retrain the model against attacks. In contrast, our method can effectively enhance the robustness through ICL without adjusting the parameters of the LLMs, which is undoubtedly a significant advantage.
5.3.2 The Challenges in LLM-based Defense
@@
The defense agent aims to purify adversarial inputs, but it is difficult to distinguish between natural examples and adversarial examples in some cases. As shown in Table 12.4, the attacker altered the original meaning by inserting ‘not’, rendering the adversarial example indistinguishable from a natural example, resulting in the defense agent failing to generate the correct sentence. Although we hope that the defense agent can observe the sentences like humans, it presents a huge challenge at present. Unlike the attacker or humans, the defense agent lacks access to the original label of the input sentence. In addition, even successful purifications can be re-attacked. As shown in Table 7, the malicious LLMs can embed specific system prompts to influence the output, which is unbeknownst to users; they can add ‘:)’ to each input sentence for prediction rather than predicting the original sentence. In this case, the defense agent also fails to defend. This issue is also an important problem in traditional adversarial training [5], and no method has completely resolved this issue. Nonetheless, as previously discussed, LLMs offer advantages not available to traditional DNNs, and we have naturally solved one challenge in adversarial robustness, which is that the model can adapt to new attacks. Hence, future advancements may resolve adversarial issues of attack and defense within LLM frameworks, representing a challenging but promising research direction.
6 Conclusion
@@
In this paper, we propose LLAMOS, a novel LLM-based defense technique designed to purify adversarial examples before feeding them into the target LLM. The defense agent within LLAMOS operates as a plug-and-play module that functions effectively as a pre-processing step without requiring retraining of the target LLM. We conduct extensive experiments across various tasks and attacks with LLAMA-2 and GPT-3.5. The results demonstrate that LLAMOS can effectively defend against adversarial attacks. Furthermore, we discuss certain existing shortcomings and challenges, which we aim to address in future research.
Appendix 0.A GLUE Dataset
0.A.0.1 Stanford Sentiment Treebank
SST-2 [28] is a single-sentence classification task, includes sentences from movie reviews and their sentiment annotations by humans. This task involves determining the sentiment of a given sentence, categorized into two types: positive and negative.
0.A.0.2 Recognizing Textual Entailment
RTE [2] is a binary classification task, where the goal is to determine whether a given sentence entails another sentence.
0.A.0.3 Quora Question Pairs
QQP [38] is a collection of question pairs from the community question-answering website Quora. The task is to determine whether a pair of questions are semantically equivalent.
0.A.0.4 Question-answering NLI
QNLI [23] is a natural language inference task, converted from another dataset, The Stanford Question Answering Dataset. The task is to determine whether a question and a sentence are entailed or not.
0.A.0.5 Multi-Genre Natural Language Inference
MNLI [39] is a collection of sentence pairs with textual entailment annotations. Given a premise sentence and a hypothesis sentence, the task is to predict whether the premise entails the hypothesis, contradicts the hypothesis, or neither. MNLI consists of texts from various domains and is divided into two versions: MNLI-m, where training and test datasets share the same sources, and MNLI-mm, where they differ.
Table 9: The task descriptions of the GLUE dataset.
Datasets [Task Description]
SST-2 Analyze the tone of this statement and respond with either ‘positive’ or ‘negative’.
RTE Are the following two sentences entailment or not_entailment? Answer me with ‘entailment’ or ‘not_entailment’, just one word.
QQP Are the following two questions equivalent or not? Answer me with ‘equivalent’ or ‘not_equivalent’.
QNLI Given the question and context provided, determine if the answer can be inferred by choosing ‘entailment’ or ‘not_entailment’.
MNLI-mm Does the relationship between the given sentences represent entailment, neutral, or contradiction? Respond with ‘entailment’, ‘neutral’, or ‘contradiction’.
MNLI-m Does the relationship between the given sentences represent entailment, neutral, or contradiction? Respond with ‘entailment’, ‘neutral’, or ‘contradiction’.
Table 10: The label list and input description (EEC = Each example contains). For instance, EEG one ‘sentence’ = Each example contains one ‘sentence’.
Datasets [Label List] [Input Description]
SST-2 [‘positive’, ‘negative’] EEC one ‘sentence’.
RTE [‘entailment’, ‘not_entailment’] EEC ‘sentence1’ and ‘sentence2’.
QQP [‘equivalent’, ‘not_equivalent’] EEC ‘question1’ and ‘question2’.
QNLI [‘entailment’, ‘not_entailment’] EEC ‘question’ and ‘sentence’.
MNLI-mm [‘entailment’, ‘neutral’, ‘contradiction’] EEC ‘premise’ and ‘hypothesis’.
MNLI-m [‘entailment’, ‘neutral’, ‘contradiction’] EEC ‘premise’ and ‘hypothesis’.
Appendix 0.B PromptAttack and More Comparasion Results
PromptAttack [41] modifies the clean examples at the character level, word level or sentence level. The specific guidance as shown in Table 11. Few-shot (FS): Provide examples that match the task description of adversarial attacks to LLMs to help the LLMs understand the task. Ensemble (EN): Utilize a collection of adversarial examples with various levels of perturbation, select the examples that are most likely to deceive the LLM as output.
Table 11: Perturbation instructions at the character, word, and sentence levels, respectively. This table is referenced from [41].
Level Abbre. #perturbation_instruction
Character C1 Choose at most two words in the sentence, and change them so that they have typos.
C2 Change at most two letters in the sentence.
C3 Add at most two extraneous characters to the end of the sentence.
Word W1 Replace at most two words in the sentence with synonyms.
W2 Choose at most two words in the sentence that do not contribute to the meaning of the sentence and delete them.
W3 Add at most two semantically neutral words to the sentence.
Sentence S1 Add a randomly generated short meaningless handle after the sentence, such as @fasuv3.
S2 Paraphrase the sentence.
S3 Change the syntactic structure of the sentence.
Table 12: Comparison between clean, adversarial, and purified examples. When incorrectly classified, the modified area is highlighted with a yellow background.
[Input Examples] Prediction
1. Clean Example: it ’s a charming and often affecting journey . Positive
Adv. Example: it ’s a charming and oftenaffecctingjourney . Negative
Purified Example: it’s a charming and often affecting journey. Positive
2. Clean Example: … the movie is just a plain old monster Negative
Adv. Example: … the movie is just a plain old monster:) Positive
Purified Example: … the movie is just a plain old monster . Negative
3. Clean Example: the year’s most intriguing explorations of alientation . Positive
Adv. Example: the year’s most intriguing explorations of alientation .:( Negative
Purified Example: the year’s most intriguing explorations of alientation.. Positive
4. Clean Example: it ’s a charming and often affecting journey . Positive
Adv. Example: it ’snota charming and often affecting journey .@fasuv3 Negative
Purified Example: it ’snota charming and often affecting journey . Negative
5. Clean Example: … to be kind of heartwarming, nonetheless. Positive
Adv. Example: … to be kind of heartwarming, nonetheless.@kjdjq2. Negative
Purified Example: … to be kind of heartwarming, nonetheless.. Positive